Using x for the variable, write an equation that describes the given situation, then solve the equation. Veronica sold her home for x dollars. The real estate agent's commission was 5% of the selling price and Veronica received $133,000.
Answer by deborabr(181)

5% = 5/100
5/100*x+133000=x
5x/100+13300000/100=100x/100
5x+13300000=100x
5x-100x=-13300000
-95x=-13300000 *(-1)
95x=13300000
x=13300000/95
x=140000